OpenStudy (anonymous):

POLICE CAN ESTIMATE THE SPEED OF A VEHICLE BEFORE THE BRAKES ARE APPLIED USING THE FORMULA , WHERE S IS THE SPEED IN MILES PER HOUR AND D IS THE LENGTH OF THE VEHICLE’S SKID MARKS. WHAT WAS THE APPROXIMATE SPEED OF A VEHICLE THAT LEFT A SKID MARK MEASURING 125 FEET?

OpenStudy (anonymous):

what formula. you asked a similar question earlier. something like .75d=s^2/2 maybe?

OpenStudy (anonymous):

using the formula \[e=mc^2\]?

OpenStudy (anonymous):

Formula = .75d=s squared/30.25

OpenStudy (anonymous):

\[\S^2=.75d(30.25)\] Only need the positive root, so \[\S= \sqrt{.75d(30.25)}\]

OpenStudy (anonymous):

what did you get as the final answer 22.7d

OpenStudy (anonymous):

They gave you d=125

OpenStudy (anonymous):

I got 2835.9

OpenStudy (anonymous):

it says round to the nearest mile per hour

OpenStudy (anonymous):

o and square root it

OpenStudy (anonymous):

Take the square root of that

OpenStudy (anonymous):

53.25

OpenStudy (anonymous):

to the nearest mph, it's 53 mph